If 190cmcube3 of hydrogen is collected at 26degree celcius and 736mmHg,calculate it volume at S.T.P (standard pressure is 760mmHg

1
Expert's answer
2022-02-01T01:34:31-0500

P1= 736

V1= 190

T1= 299.15

T2 = 273.15

P2= 760

V2= ?

"P1v1T2= P2V2T1"

736× 190× 273.15= 760× 299.15× V2

(41833136/227354)=

V2= 184cm3